OpenStudy (anonymous):

rewrite the equation in slope intersept form then identify the slope and y intercept of the line. x-y=6

OpenStudy (anonymous):

y = x - 6 the slope is 1. When x=0, y=-6. Remember: y -y0 = m(x-x0)

OpenStudy (anonymous):

oh ok thanks

OpenStudy (anonymous):

you're welcome!